A child weighs 85lbs and is 50 inches tall. The recommended dose of vincristine is 1.5-2mg/m2 once a week. The concentration available is 1mg/mL. What is the recommended dosage range for this child?
Answer #1

ANSWER.

* Calculation of recommended dose for this child.

      Recommended dose = 1.5 - 2 mg/m2.

         Body surface area (BSA) m2.

           Formula

                    m2 = square root (height * weight ) / 3600.

            

        Weight of the patient = 85 lbs.

       Convert lbs to kg for calculation.

                   1 lbs = 0.45 kg.

        So

            Weight of the patient = 85 * 0.454

                                              = 38.6 kg.

           Height of the patient = 50 inch.

        Convert inch to cm for calculation.

                      1 inch = 2.54 cm.

        So

            Height of the patient = 50 * 2.54

                                             = 127 cm.

      Body surface area of this child = Square root (127 * 38.6) / 3600

                                                             = Square root (4902.2) / 3600

                                                             = Square root (1.36)

                                                             = 1.17 m2.

          Recommended dose for this child = 1.5 * 1.17 - 2 * 1.17

                                                           = 1.8mg - 2.3mg once a week.